QUESTION 15 When I get my tax refund this year, theres a 30% chance l will buy a wide screen TV a 20% chance I will buy a blue ray player, and a 10% chance I will buy both. Whats the chance will buy neither and hit Foxwoods instead? (Hint: Complement of PCTV or Blue Ray)) OA.9496 OB, 4096 О с. 60% on 99.4% QUESTION 16 Suppose 1 in 5 (1/5) JwU students own their own car. If four students are randomly selected, what is the probability that all four own their car? O A. 1/625 O B. 4/20 O C. 4/5 O D. 1/20
Answer #1

15.

Solution:-

P(TV)=30%, P(Player)=20%, P(TV \cap Player)=10%

P(TV U Player)=P(TV)+ P(Player)- P(TV \cap Player)

=30+20-10

=40%

Probability that I will buy Neither=P(TV U Player)'

=100%-P(TV U Player)

=100%-40%

=60% ...ans=C.